# Written by AAAI Press Staff1

### v1

@@ Abstract
Graph partitioning aims to divide a graph into $k$ disjoint subsets while optimizing a specific partitioning objective. The majority of formulations related to graph partitioning exhibit NP-hardness due to their combinatorial nature. As a result, conventional approximation algorithms rely on heuristic methods, sometimes with approximation guarantees and sometimes without. Unfortunately, traditional approaches are tailored for specific partitioning objectives and do not generalize well across other known partitioning objectives from the literature. To overcome this limitation, and learn heuristics from the data directly, neural approaches have emerged, demonstrating promising outcomes. In this study, we extend this line of work through a novel framework, NeuroCUT. NeuroCUT introduces two key innovations over prevailing methodologies. First, it is inductive to both graph topology and the partition count, which is provided at query time. Second, by leveraging a reinforcement learning based framework over node representations derived from a graph neural network, NeuroCUT can accommodate any optimization objective, even those encompassing non-differentiable functions. Through empirical evaluation, we demonstrate that NeuroCUT excels in identifying high-quality partitions, showcases strong generalization across a wide spectrum of partitioning objectives, and exhibits resilience to topological modifications.
@@ 1 Introduction and Related Work
Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (karypis1997multilevel; karypis1998multilevelk; karypis1999multilevel; andersen2006local; chung2007four). The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (karypis1997multilevel).
@@
The most relevant method to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap). It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation. Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (duong2023deep). Another method solves the multicut problem where the number of partitions is not an input to the problem (jung2022learning). The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function. Finally,  (gatti2022graph) solves the normalized cut problem only for the case where the number of partitions is exactly two.
Nevertheless, existing neural approaches that tackle graph partitioning problem suffer three notable drawbacks, which we address.
  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable. As we will illustrate in § 2, the assumption does not always hold in the real-world.
@@
  • Inductivity to partition count: The number of partitions required to segment a graph is an input parameter. Hence, it is important for a learned model to generalize to any partition count without retraining. Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained. Additionally, it’s worth noting that the optimal number of partitions is often unknown beforehand. Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective. Typically, the optimal partition count corresponds to the point where the second-order derivative of the objective function approaches saturation. Therefore, inductivity to the number of partitions becomes an essential requirement for conducting such experiments in a scalable manner.
  • Generalizability to objective functions: Multiple objective functions have been studied in the partitioning literature. The optimal objective function hinges upon the subsequent application in question. Consequently, it holds significance for an algorithmic framework to possess the capacity to encompass a diverse array of objective functions. Unfortunately, prevailing neural frameworks are founded upon presumptions that yield efficacy exclusively within the confines of their designated objective function.
@@ 1.1 Contributions:
  • Versatile objectives: We develop an auto-regressive, graph reinforcement learning framework, called NeuroCUT, to solve the graph partitioning problem. Diverging from conventional algorithms, NeuroCUT can handle different partitioning objectives. Moreover, unlike other neural methods, NeuroCUT can accommodate diverse partitioning objectives, without the necessity for differentiability.
  • Inductivity to topology and number of partitions: The parameter space of NeuroCUT is independent of the graph size or the partition count. This innovative architecture endows NeuroCUT with the ability to generalize effectively to previously unseen graphs of varying scales and adapt to arbitrary partition count specified during inference.
  • Empirical Assessment: We perform comprehensive experiments employing real-world datasets, evaluating NeuroCUT across four distinct graph partitioning objectives. Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions, and demonstrating its capacity to generalize effectively to partition sizes and topological configurations that it has not encountered during training.
@@ 2 Problem Formulation
In this section, we introduce the concepts central to our work and formulate the problem. All the notations used in this work are outlined in Tab. 1 in appendix.
@@
We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{G}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges.
@@
    Normalized Mincut (shi2000normalized): The $k$-minimum cut criteria favors cutting small sets of isolated nodes in the graph. To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
@@
    (Weighted) Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition. We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions. The intuition behind sparsest cuts is that any partition should neither be very large nor very small.
@@
Figure 1: Architecture of NeuroCUT. First the initial partitioning of the graph is performed based upon node’s initial embeddings. The initial embeddings are refined using GNN to infuse toplogical information from neighborhood. At each step a node is selected and its partitioned is updated. During the training phase the GNN parameters are updated and hence embeddings are re-computed. During inference, the GNN is called only once to compute the embeddings of the nodes of the graph.
@@
Initial feature representation of node $v$
@@ 3 NeuroCUT: Proposed Methodology
Fig. 1 describes the framework of NeuroCUT. For a given input graph $\mathcal{G}$, we first construct the initial partitions of nodes using a clustering based approach. Subsequently, a message-passing Gnn embeds the nodes of the graph ensuring inductivity to different graph sizes. Next, the assignment of nodes to partitions proceeds in a two-phased strategy. We first select a node to change its partition and then we choose a suitable partition for the selected node. Further, to ensure inductivity on the number of partitions, we decouple the number of partitions from the direct output representations of the model. This decoupling allows us to query the model to unseen number of partitions. After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized. The reward us learned through reinforcement learning (RL) (sutton2018reinforcement).
@@
The choice of using RL is motivated two observations. Firstly, cut problems on graphs are generally recognized as NP-hard, making it impractical to rely on ground-truth data, which would be computationally infeasible to obtain. Secondly, the cut objective may lack differentiability. Therefore, it becomes essential to adopt a learning paradigm that can be trained even under these non-differentiable constraints. In this context, RL effectively addresses both of these critical requirements.
We now discuss each of the components in detail.
3.1 Learning Policy $\pi$ as Markov Decision Process
@@
Given a graph $\mathcal{G}$, our objective is to find the partitioning $\mathcal{P}$ that maximizes/minimizes the target objective function $Obj(\mathcal{G},\mathcal{P})$. We model the task of iteratively updating the partition for a node as a Markov Decision Process (MDP) defined by the tuple $(S,\mathcal{A},\rho,R,\gamma)$. Here, $S$ is the state space, $\mathcal{A}$ is the set of all possible actions, $\rho:S\times S\times\mathcal{A}\rightarrow[0,1]$ denotes the state transition probability function, $R:S\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function and $\gamma\in(0,1)$ the discounting factor. We next formalize each of these notions in our MDP formulation.
@@
Initialization. Instead of directly starting from empty partitions, we perform a warm start operation that clusters the nodes of the graph to obtain the initial graph partitions. The graph is clustered into $k$ clusters based upon their positional embeddings (discussed below). We use K-means (macqueen1967some) algorithm for this task. The details of the clustering are present in Appendix B.
@@
Positional Embeddings: Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph. Positional encodings provide an idea of the position in space of a given node within the graph. Two nodes that are closer in the graph, should be closer in the embedding space. Towards this, we use Lipschitz Embedding (bourgain1985lipschitz). Let $\mathcal{A}=\{a_{1},\cdots,a_{\alpha}\}\subseteq\mathcal{V}$ be a randomly selected subset of $\alpha$ nodes. We call them anchor nodes. From each anchor node $i$, the walker starts a random walk (bianchini2005inside) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$. Furthermore, at each step, with probability $c$ the walker jumps to a neighboring node $j$ and returns to the node $i$ with $1-c$. Let $\vec{r}_{ij}$ corresponds to the probability of the random walker starting from node $i$ and reaching node $j$.
@@
State: The state space characterizes the state of the system at time $t$ in terms of the current set of partitions $\mathcal{P}^{t}$. Intuitively the state should contain information to help our model make a decision to select the next node and the partition for the node to be assigned. Let $\mathcal{P}^{t}$ denote the status of partitions at time $t$ wherein a partition $P^{t}_{i}$ is represented by all nodes belonging to $i^{th}$ partition. The state of the system at step $t$ is defined as
$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{pos}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (8)
Here state of each partition is represented by the collection of positional embedding of nodes in that partition.
Action: Towards finding the best partitioning scheme for the target partition objective we propose a 2-step action strategy to update the partitions of nodes. The first phase consists of identifying a node to update its partition. Instead of arbitrarily picking a node, we propose to prioritize selecting nodes for which the new assignment is more likely to improve the overall partitioning objective. In comparison to a strategy that arbitrarily selects nodes, the above mechanism promises greater improvement in the objective with less number of iterations. In the second phase, we calculate the score of each partition $\mathcal{P}$ with respect to the selected node from the first phase and then assign it to one of the partitions based upon the partitioning scores. We discuss both these phases in details below.
Phase 1: Node Selection Let $\operatorname{PART}(\mathcal{P}^{t},v)$ denote the partition of the node $v$ at step $t$. Our proposed formulation involves selecting a node $v$ at step $t$ belonging to partition $\operatorname{PART}(\mathcal{P}^{t},v)$ and then assigning it to a new partition. The newly assigned partition and the current partition of the node could also be same.
@@
$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in\mathcal{P}^{t}-\operatorname{PART}(\mathcal{P}^{t},v)}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=p|}}}{|\forall u\in\mathcal{N}(v)\ni\operatorname{PART}(\mathcal{P}^{t},u)=\operatorname{PART}(\mathcal{P}^{t},v)|}}$ (9)
@@
Phase 2: Inductive Neural Method for Partition Selection Once a node is selected, the next phase involves choosing the new partition for the node. Towards this, we design an approach empowered by Graph Neural Networks (GNNs) which enables the model to be inductive with respect to size of graph. Further, instead of predicting a fixed-size score vector (nazi2019gap) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too. We discuss both above points in section below.
Message passing through graph neural network: To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (gcn; velivckovic2017graph; hamilton2017inductive). GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{pos}(u)$ using eq. 6. We perform $L$ layers of message passing to compute representations of nodes. To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(hamilton2017inductive):
$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\mathrm{mean}_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (10)
@@
Scoring partitions: Recall from eq. 8, each partition at time $t$ is represented using the nodes belonging to that partition. Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$. In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap), the proposed design choice makes the model inductive to the number of partitions. Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
@@
$\pi((a^{t}{=}p)/S^{t})=\frac{exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (12)
@@
State Transition: After action $a^{t}$ is applied at state $S^{t}$, the state is updated to $S^{t+1}$ that involves updating the partition set $\mathcal{P}^{t+1}$. Specifically, if node $v$ belonged to $i^{th}$ partition at time $t$ and its partition has been changed to $j$ in phase 2, then we apply the below operations in order.
@@
Reward: Our aim is to improve the value of the overall partitioning objective. One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$, i.e., $R^{t}=(Obj(\mathcal{G},\mathcal{P}^{t+1})-Obj(\mathcal{G},\mathcal{P}^{t}))/(Obj(\mathcal{G},\mathcal{P}^{t+1})+Obj(\mathcal{G},\mathcal{P}^{t}))$. Here, the denominator term ensures that the model receives a significant reward for even slight enhancements when dealing with a small objective value. However, this definition of reward focuses on short-term improvements instead of long-term. Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term. The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term). Mathematically,
$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}\vspace{-0.1in}$ (14)

### v2

@@ Abstract
Graph partitioning aims to divide a graph into $k$ disjoint subsets while optimizing a specific partitioning objective. The majority of formulations related to graph partitioning exhibit NP-hardness due to their combinatorial nature. Conventional methods, like approximation algorithms or heuristics, are designed for distinct partitioning objectives and fail to achieve generalization across other important partitioning objectives. Recently machine learning-based methods have been developed that learn directly from data. Further, these methods have a distinct advantage of utilizing node features that carry additional information. However, these methods assume differentiability of target partitioning objective functions and cannot generalize for an unknown number of partitions, i.e., they assume the number of partitions is provided in advance. In this study, we develop NeuroCUT with two key innovations over previous methodologies. First, by leveraging a reinforcement learning-based framework over node representations derived from a graph neural network and positional features, NeuroCUT can accommodate any optimization objective, even those with non-differentiable functions. Second, we decouple the parameter space and the partition count making NeuroCUT inductive to any unseen number of partition, which is provided at query time. Through empirical evaluation, we demonstrate that NeuroCUT excels in identifying high-quality partitions, showcases strong generalization across a wide spectrum of partitioning objectives, and exhibits strong generalization to unseen partition count.
@@ 1 Introduction and Related Work
Several graph partitioning formulations have been studied in the literature, mostly in the form of discrete optimization (karypis1997multilevel; karypis1998multilevelk; karypis1999multilevel; andersen2006local; chung2007four). The majority of the formulations are NP-hard and thus the proposed solutions are either heuristics or algorithms with approximate solutions (karypis1997multilevel). Recently, there have been attempts to solve graph partitioning problem via neural approaches. The neural approaches have a distinct advantage that they can utilize node features. Node features supply additional information and provide contextual insights that may improve the accuracy of graph partitioning. For instance, (wang2017mgae) has recognized the importance of incorporating node attributes in graph clustering where nodes are partitioned into disjoint groups. Note that there are existing neural approaches (khalil2017learning; gcomb; kool2018attention; joshi2019efficient) to solve other NP-hard graph combinatorial problems (e.g., minimum vertex cover). However these methods are not generic enough to solve graph partitioning.
@@
In this paper, we build a single framework to solve several graph partitioning problems. One of the most relevant to our work is the method DMoN by muller2023graph. This method designs a neural architecture for cluster assignments and use a modularity-based objective function for optimizing these assignments. Another method, that is relevant to our work is GAP, which is an unsupervised learning method to solve the balanced graph partitioning problem (nazi2019gap). It proposes a differentiable loss function for partitioning based on a continuous relaxation of the normalized cut formulation. Deep-MinCut being an unsupervised approach learns both node embeddings and the community structures simultaneously where the objective is to minimize the mincut loss (duong2023deep). Another method solves the multicut problem where the number of partitions is not an input to the problem (jung2022learning). The idea is to construct a reformulation of the multicut ILP constraints to a polynomial program as a loss function. However, the problem formulation is different than the generic normalized cut or mincut problem. Finally,  (gatti2022graph) solves the normalized cut problem only for the case where the number of partitions is exactly two. Nevertheless, these neural approaches for the graph partitioning problem often do not use node features and only limited to a distinct partitioning objective. Here, we point out notable drawbacks that we address in our framework.
@@
  • Non-inductivity to partition count: The number of partitions required to segment a graph is an input parameter. Hence, it is important for a learned model to generalize to any partition count without retraining. Existing neural approaches are non-inductive to the number of partitions, i.e they can only infer on number of partitions on which they are trained. Additionally, it is worth noting that the optimal number of partitions is often unknown beforehand. Hence, it is a common practice to experiment with different partition counts and evaluate their impact on the partitioning objective. While the existing methods (muller2023graph; nazi2019gap; gatti2022graph) assume that the number of partitions ($k$) is known beforehand, our proposed method can generalize to any $k$.
  • Non-generalizability to different cut functions: Multiple objective functions for graph cut have been studied in the partitioning literature. The optimal objective function hinges upon the subsequent application in question. For instance, the two most relevant studies, DMoN (muller2023graph) and GAP (nazi2019gap) focus on maximizing modularity and minimizing normalized cut respectively. Our framework is generic and can solve different partitioning objectives.
  • Assumption of differential objective function: Existing neural approaches assume the objective function to be differentiable. As we illustrate in § 2, the assumption does not always hold in the real-world. For instance, the sparsest cut (chawla2006hardness) and balanced cut (nazi2019gap) objectives are not differentiable.
@@ 1.1 Contributions:
  • Versatile objectives: We develop NeuroCUT; an auto-regressive, graph reinforcement learning framework that integrates positional information, to solve the graph partitioning problem for attributed graphs. Diverging from conventional algorithms, NeuroCUT can solve multiple partitioning objectives. Moreover, unlike other neural methods, NeuroCUT can accommodate diverse partitioning objectives, without the necessity for differentiability.
  • Inductivity to number of partitions: The parameter space of NeuroCUT is independent of the partition count. This innovative decoupled architecture endows NeuroCUT with the ability to generalize effectively to arbitrary partition count specified during inference.
  • Empirical Assessment: We perform comprehensive experiments employing real-world datasets, evaluating NeuroCUT across four distinct graph partitioning objectives. Our empirical investigation substantiates the efficacy of NeuroCUT in partitioning tasks, showcasing its robustness across a spectrum of objective functions. We also demonstrate the capacity of NeuroCUT to generalize effectively to partition sizes that it has not encountered during training.
@@ 2 Problem Formulation
In this section, we introduce the concepts central to our work and formulate the problem. All the notations used in this work are outlined in Table 1.
@@
We denote a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E},\textbf{{X}})$ where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges and $\textbf{{X}}\in\mathbb{R}^{|\mathcal{V}|\times|F|}$ refers to node feature matrix where $F$ is the set of all features in graph $\mathcal{G}$.
@@
    Normalized Cut (shi2000normalized): The $k$-MinCut criteria favors cutting small sets of isolated nodes in the graph. To avoid this unnatural bias for partitioning out small sets of points, normalised cut computes the cut cost as a fraction of the total edge connections to all the nodes in the graph.
@@
    Sparsest Cuts (chawla2006hardness): Two-way sparsest cuts minimize the cut edges relative to the number of nodes in the smaller partition. We generalize it to $k$-way sparsest cuts by summing up the value for all the partitions. The intuition behind sparsest cuts is that any partition should neither be very large nor very small.
@@
Figure 1: Architecture of NeuroCUT. First, the initial partitioning of the graph is performed based on node raw features and positional embeddings. These embeddings are refined using GNN to infuse toplogical information from neighborhood. At each step a node is selected and its partitioned is updated. During the training phase the GNN parameters are updated and hence embeddings are re-computed. During inference, the GNN is called only once to compute the embeddings of the nodes of the graph.
@@
Positional embedding of node $v$
$\mathbf{emb_{init}}(v)$
Initial embedding of node $v$
@@ 3 NeuroCUT: Proposed Methodology
Fig. 1 describes the framework of NeuroCUT. For a given input graph $\mathcal{G}$, we first construct the initial partitions of nodes using a clustering based approach. Subsequently, a message-passing Gnn embeds the nodes of the graph ensuring inductivity to different graph sizes. Next, the assignment of nodes to partitions proceeds in a two-phased strategy. We first select a node to change its partition and then we choose a suitable partition for the selected node. Further, to ensure inductivity on the number of partitions, we decouple the number of partitions from the direct output representations of the model. This decoupling allows us to query the model to an unseen number of partitions. After a node’s partition is updated, the reward with respect to change in partitioning objective value is computed and the parameters of the policy are optimized. The reward is learned through reinforcement learning (RL) (sutton2018reinforcement).
@@
The choice of using RL is motivated through two observations. Firstly, cut problems on graphs are generally recognized as NP-hard, making it impractical to rely on ground-truth data, which would be computationally infeasible to obtain. Secondly, the cut objective may lack differentiability. Therefore, it becomes essential to adopt a learning paradigm that can be trained even under these non-differentiable constraints. In this context, RL effectively addresses both of these critical requirements. Additionally, in the process of sequentially constructing a solution, RL allows us to model the gain obtained by perturbing the partition of a node.
@@
Markov Decision Process. Given a graph $\mathcal{G}$, our objective is to find the partitioning $\mathcal{P}$ that maximizes/minimizes the target objective function $Obj(\mathcal{G},\mathcal{P})$. We model the task of iteratively updating the partition for a node as a Markov Decision Process (MDP) defined by the tuple $(S,\mathcal{A},\rho,R,\gamma)$. Here, $S$ is the state space, $\mathcal{A}$ is the set of all possible actions, $\rho:S\times S\times\mathcal{A}\rightarrow[0,1]$ denotes the state transition probability function, $R:S\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function and $\gamma\in(0,1)$ the discounting factor. We next formalize each of these notions in our MDP formulation.
3.1 State: Initialization & Positional Encoding
@@
Initialization. Instead of directly starting from empty partitions, we perform a warm start operation that clusters the nodes of the graph to obtain the initial graph partitions. The graph is clustered into $k$ clusters based on their raw features and positional embeddings (discussed below). We use K-means (macqueen1967some) algorithm for this task. The details of the clustering are present in Appendix A.2.
@@
Positional Encoding (Embeddings). Given that the partitioning objectives are NP-hard mainly due to the combinatorial nature of the graph structure, we look for representations that capture the location of a node in the graph. Positional encodings provide an idea of the position in space of a given node within the graph. Two nodes that are closer in the graph, should be closer in the embedding space. Towards this, we use Lipschitz Embedding (bourgain1985lipschitz). Let $\mathcal{A}=\{a_{1},\cdots,a_{\alpha}\}\subseteq\mathcal{V}$ be a randomly selected subset of $\alpha$ nodes. We call them anchor nodes. From each anchor node $i$, the walker starts a random walk (bianchini2005inside) and jumps to a neighboring node $j$ with a transition probability ($\mathbf{W}_{ij}$) governed by the transition probability matrix $\mathbf{W}\in\mathbb{R}^{|V|\times|V|}$. Furthermore, at each step, with probability $c$ the walker jumps to a neighboring node $j$ and returns to the node $i$ with $1-c$. Let $\vec{r}_{ij}$ corresponds to the probability of the random walker starting from node $i$ and reaching node $j$.
@@
To accommodate both raw feature and positional information we concatenate the raw node features i.e $\textbf{X}[u]$ with the positional embedding $\mathbf{pos}(u)$ for each node $u$ to obtain the initial embedding which will act as input to our neural model. Specifically,
$\mathbf{emb_{init}}(u)=\textbf{X}[u]\;\;\mathbin{\|}\;\;\mathbf{pos}(u)$ (8)
[…]
State. The state space characterizes the state of the system at time $t$ in terms of the current set of partitions $\mathcal{P}^{t}$. Intuitively the state should contain information to help our model make a decision to select the next node and the partition for the node to be assigned. Let $\mathcal{P}^{t}$ denote the status of partitions at time $t$ wherein a partition $P^{t}_{i}$ is represented by all nodes belonging to $i^{th}$ partition. The state of the system at step $t$ is defined as
$S^{t}=\{S_{1}^{t},S_{2}^{t},\cdots S_{k}^{t}:\;S_{i}^{t}{=}\{\mathbf{emb_{init}}(v)\;\;\forall v\in P_{i}^{t}\}\}$ (9)
Here state of each partition is represented by the collection of initial embedding of nodes in that partition.
3.2 Action: Selection of Nodes and Partitions
Towards finding the best partitioning scheme for the target partition objective we propose a 2-step action strategy to update the partitions of nodes. The first phase consists of identifying a node to update its partition. Instead of arbitrarily picking a node, we propose to prioritize selecting nodes for which the new assignment is more likely to improve the overall partitioning objective. In comparison to a strategy that arbitrarily selects nodes, the above mechanism promises greater improvement in the objective with less number of iterations. In the second phase, we calculate the score of each partition $\mathcal{P}$ with respect to the selected node from the first phase and then assign it to one of the partitions based upon the partitioning scores. We discuss both these phases in details below.
Phase 1: Node selection to identify node to perturb
Let $\operatorname{PART}(\mathcal{P}^{t},v)$ denote the partition of the node $v$ at step $t$. Our proposed formulation involves selecting a node $v$ at step $t$ belonging to partition $\operatorname{PART}(\mathcal{P}^{t},v)$ and then assigning it to a new partition. The newly assigned partition and the current partition of the node could also be same.
@@ 3.2 Action: Selection of Nodes and Partitions
$\operatorname{NodeScore}^{t}[v]=\\ {\frac{{\max_{p\in P^{t}-\operatorname{PART}(P^{t},v)}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=p|}}}{|u|u\in\mathcal{N}(v)\ni\operatorname{PART}(P^{t},u)=\operatorname{PART}(P^{t},v)|}}$ (10)
@@
Phase 2: Inductive Method for Partition Selection
Once a node is selected, the next phase involves choosing the new partition for the node. Towards this, we design an approach empowered by Graph Neural Networks (GNNs) (hamilton2017inductive) which enables the model to be inductive with respect to size of graph. Further, instead of predicting a fixed-size score vector (nazi2019gap; tsitsulin2023graph) for the number of partitions, our proposed method of computing partition scores allows the model to be inductive to the number of partitions too. We discuss both above points in section below.
Message passing through Graph Neural Network
To capture the interaction between different nodes and their features along with the graph topology, we parameterize our policy by a Graph Neural Network (GNN) (hamilton2017inductive). GNNs combine node feature information and the graph structure to learn better representations via feature propagation and aggregation.
We first initialize the input layer of each node $u\in\mathcal{V}$ in graph as $\mathbf{h}^{0}_{u}=\mathbf{emb_{init}}(u)$ using eq. 8. We perform $L$ layers of message passing to compute representations of nodes. To generate the embedding for node $u$ at layer $l+1$ we perform the following transformation(hamilton2017inductive):
$\mathbf{h}_{u}^{l+1}$ $=\mathbf{W}^{l}_{1}\mathbf{h}_{u}^{l}+\mathbf{W}^{l}_{2}\cdot\frac{1}{|\mathcal{N}_{u}|}\sum_{u^{\prime}\in\mathcal{N}_{u}}\mathbf{h}_{u^{\prime}}^{l}$ (11)
@@
Scoring partitions. Recall from eq. 9, each partition at time $t$ is represented using the nodes belonging to that partition. Building upon this, we compute the score of each partition $p\in\mathcal{P}^{t}$ with respect to the node $v$ selected in Phase 1 using all the nodes in $p$. In contrast to predicting a fixed-size score vector corresponding to number of partitions (nazi2019gap; tsitsulin2023graph), the proposed design choice makes the model inductive to the number of partitions. Specifically, the number of partitions are not directly tied to the output dimensions of the neural model.
@@
$\pi((a^{t}{=}p)/S^{t})=\frac{\exp\left(\operatorname{PartScore}\left(p,v\right)\right)}{\sum_{p^{\prime}\in\mathcal{P}^{t}}\exp\left(\operatorname{PartScore}\left(p^{\prime},v\right)\right)}$ (13)
@@
State Transition. After action $a^{t}$ is applied at state $S^{t}$, the state is updated to $S^{t+1}$ that involves updating the partition set $\mathcal{P}^{t+1}$. Specifically, if node $v$ belonged to $i^{th}$ partition at time $t$ and its partition has been changed to $j$ in phase 2, then we apply the below operations in order.
@@
3.3 Reward, Training & Inference
Reward. Our aim is to improve the value of the overall partitioning objective. One way is to define the reward $R^{t}$ at step $t\geq 0$ as the change in objective value of the partitioning i.e $Obj(\mathcal{G},\mathcal{P}^{t})$ at step $t$, i.e., $R^{t}=(Obj(\mathcal{G},\mathcal{P}^{t+1})-Obj(\mathcal{G},\mathcal{P}^{t}))/(Obj(\mathcal{G},\mathcal{P}^{t+1})+Obj(\mathcal{G},\mathcal{P}^{t}))$. Here, the denominator term ensures that the model receives a significant reward for even slight enhancements when dealing with a small objective value. However, this definition of reward focuses on short-term improvements instead of long-term. Hence, to prevent this local greedy behavior and to capture the combinatorial aspect of the selections, we use discounted rewards $D^{t}$ to increase the probability of actions that lead to higher rewards in the long term (sutton2018reinforcement). The discounted rewards are computed as the sum of the rewards over a horizon of actions with varying degrees of importance (short-term and long-term). Mathematically,
$D^{t}=R^{t}+\gamma R^{t+1}+\gamma^{2}R^{t+2}+\ldots=\sum_{j=0}^{T-t}\gamma^{j}R^{t+j}$ (15)